OpenStudy (brightemmerald):

Raffle tickets are being sold for a fundraiser. The relationship between n, number of raffle tickets sold, and A, amount of money in dollars raised, is given by the function below. A(n) = 4n - 25 What is the equation of the inverse function N(a)?

OpenStudy (brightemmerald):

Choices are; N(a)= (a+25)/4 N(a)- (a-25)/4 N(a)= a/4 -25 N(a)= a/4 +25

OpenStudy (steve816):

To find the inverse, just solve for n!